# Decoding Time Series with LLMs: A Multi-Agent Framework for Cross-Domain Annotation

### v1

@@ Abstract
Time series data is ubiquitous across various domains, including manufacturing, finance, and healthcare. High-quality annotations are essential for effectively understanding time series and facilitating downstream tasks; however, obtaining such annotations is challenging, particularly in mission-critical domains. In this paper, we propose TESSA, a multi-agent system designed to automatically generate both general and domain-specific annotations for time series data. TESSA introduces two agents: a general annotation agent and a domain-specific annotation agent. The general agent captures common patterns and knowledge across multiple source domains, leveraging both time-series-wise and text-wise features to generate general annotations. Meanwhile, the domain-specific agent utilizes limited annotations from the target domain to learn domain-specific terminology and generate targeted annotations. Extensive experiments on multiple synthetic and real-world datasets demonstrate that TESSA effectively generates high-quality annotations, outperforming existing methods.
@@ 1 Introduction
Time series data is prevalent in fields such as manufacturing Hsu and Liu (2021), finance Lee et al. (2024), and healthcare Cascella et al. (2023), where it captures critical temporal patterns essential for informed decision-making. However, general users frequently encounter difficulties in interpreting this data due to its inherent complexity, particularly in multivariate contexts where multiple variables interact over time. Furthermore, effective interpretation typically requires domain-specific knowledge to properly contextualize these patterns, thereby posing significant challenges for individuals without specialized expertise.
@@
High-quality annotations are crucial for addressing these interpretive challenges. Annotations provide meaningful context or insights into the time series data, highlighting important patterns, events, or anomalies. They facilitate accurate analysis, forecasting, and decision-making, enhancing the performance of downstream tasks such as anomaly detection, trend prediction, and automated reporting. For instance, in predictive maintenance, understanding sensor data trends is vital for preventing equipment failure, while in finance, interpreting stock price movements is crucial for informed investment strategies. Despite their importance, high-quality annotations are often scarce in real-world applications. This scarcity stems primarily from the reliance on domain experts for manual annotation, which is resource-intensive, costly, and prone to inconsistencies. Moreover, the need for precise and domain-specific terminology further complicates the annotation process, as different fields require highly specialized knowledge for accurate and contextually relevant interpretation.
@@
To alleviate the above issues, one straightforward approach is to leverage external resources to generate annotations Liu et al. (2024a). For example, Time-MMD Liu et al. (2024a) uses web searches to retrieve information as annotations, aiming to find similar patterns and descriptions from the internet. Others Jin et al. (2024); Liu et al. (2024b) directly apply large language models (LLMs) for annotation, leveraging their vast language understanding capabilities. Prototype-based methods, such as prototype networks Ni et al. (2021), have also been employed to identify representative examples for annotation. However, these methods often fall short of producing high-quality annotations. Web search-based methods may retrieve irrelevant or inconsistent information. LLMs, while powerful, tend to generate annotations that are generic, capture only basic patterns, or even hallucinate, failing to account for the complex nature of time series data. Prototype networks rely on large amounts of data to train the network and identify representative prototypes, but the scarcity of high-quality annotations limits the quality and representativeness of these prototypes, making it difficult to generalize effectively to new or unseen patterns.
@@
To address these limitations, we propose to extract knowledge from existing annotations across multiple source domains and transfer this knowledge to target domains with limited annotations. Specifically, we aim to develop a system that can automatically interpret time series data across various fields using either common or domain-specific language. Formally, given abundant annotations from multiple source domains and limited annotations from a target domain, our goal is to leverage both time-series-wise and text-wise knowledge to generate accurate and contextually appropriate annotations for the target domain.
@@
There are two major technical challenges in developing such a system: (i) How to extract common knowledge from source domains? (ii) How to learn domain-specific jargon from limited target-domain annotations? To tackle these challenges and overcome the limitations of existing methods, we propose TESSA, a multi-agent system designed for both general and domain-specific TimE SerieS Annotation. As illustrated in Figure 1, TESSA introduces two agents: a general annotation agent and a domain-specific annotation agent. The general annotation agent focuses on capturing common patterns and knowledge across various domains to generate annotations understandable by general users. To learn common knowledge from multiple domains, the general agent employs a time series-wise feature extractor and a text-wise feature extractor to extract both time-series-wise and text-wise features from time series data and domain-specific annotations from multiple source domains. To ensure important features are included in the general annotations, two feature selection methods—LLM-based and reinforcement learning-based selection—are introduced to effectively and efficiently select both the top-$k$ most important time-series-wise and text-wise features. The domain-specific agent leverages limited target-domain annotations to learn and generate annotations for specific domains using domain-specific terminologies (jargon). It incorporates a domain-specific term extractor to learn jargon from the limited target-domain annotations. Additionally, an annotation reviewer is proposed to maintain consistency between general annotations and domain-specific annotations.
@@
Our contributions are: (i) Problem. We explore a novel problem in cross-domain multi-modal time series annotation, bridging the gap between general understanding and domain-specific interpretation; (ii) Framework. We propose a novel multi-agent system, TESSA, designed for both general and domain-specific time series annotation by leveraging both time-series-wise and text-wise knowledge from multiple domains; (iii) Datasets. We collect a real-world dataset from finance domain to leverage cross-domain knowledge, along with a synthetic dataset to evaluate TESSA. These datasets are released to support future research and development in this field. (iv) Experiments. Extensive experiments on multiple synthetic and real-world datasets demonstrate the quality of the general and domain-specific annotations generated by TESSA.
@@ 2 Related Work
LLMs for Time Series Analysis. Recent advancements in LLMs have showcased their strong capabilities in sequential modeling and pattern recognition, opening up promising new directions for time series analysis. Several studies Xue and Salim (2023); Yu et al. (2023); Gruver et al. (2024); Jin et al. (2024); Li et al. (2024) have explored how LLMs can be effectively leveraged in this context. For instance, PromptCast Xue and Salim (2023) is a pioneering work that applies LLMs to general time series forecasting using a sentence-to-sentence approach. Yu et al. Yu et al. (2023) extend this by investigating the application of LLMs to domain-specific tasks, such as financial time series forecasting. LLMTime Gruver et al. (2024) demonstrates the efficacy of LLMs as time series learners by employing text-wise tokenization to represent time series data. Time-LLM Jin et al. (2024) reprograms time series data into textual prototypes for input into LLaMA-7B, enriched with natural language prompts that include domain expert knowledge and task-specific instructions. Additionally, Li et al. Li et al. (2024) illustrate how a frozen language model can enhance zero-shot learning in ECG time series analysis, showing the potential of LLMs to extract valuable features from complex time series data.
@@ 3.1 Multi-modal Feature Extraction
Time Series Feature Extraction. We develop a time series extraction toolbox $f_{t}$ to extract various features from given time series data. Formally, for each channel $c\in C$, the set of text-series features $\mathbf{F}_{t}$ is denoted as:
@@
where $f^{i}_{t}$ is the $i$-th extracted feature of $\mathbf{X}$, and $n_{t}$ is the number of extracted features. For multivariate time series data, inter-variable features (e.g., Pearson correlation) are also included. More details on feature extraction can be found in Appendix C.
@@
To address these challenges and facilitate knowledge transfer from source to target domains, we introduce a domain decontextualization LLM to convert domain-specific annotations into general annotations by removing domain-specific terminology. This makes it easier to extract common knowledge across domains. Specifically, given a domain-specific annotation $e_{s}$ in domain $d_{i}$, the decontextualized annotation $e_{d}$ is obtained as:
$e_{d}=\mathcal{M}_{d}(p_{de}(e_{s},d_{i})),$ (2)
@@
where $p_{l}$ is the prompt for text feature extraction.
@@ 3.2 Adaptive Feature Selection
With a diverse set of features extracted from time series and text data, it becomes essential to focus on the most relevant ones to ensure the generated annotations remain concise and interpretable. Moreover, repeatedly querying LLMs with both the old and new data each time wastes computational resources and incurs additional costs, especially when using non-open-source models.
@@
Offline LLM-based Feature Selection. Leveraging LLMs’ reasoning abilities, we introduce a feature selection method using LLM-generated feature importance scores to identify the top-$k$ most important text-series-wise and text-wise features. Features mentioned more frequently—either explicitly or implicitly—in annotations are assigned higher importance scores.
@@
Here, $p_{score}$ is the prompt used to score feature importance. Higher scores, $s_{j}$ and $s_{k}\in\mathbb{R}^{+}$, indicate that the features $f_{t}^{j}$ and $f_{l}^{k}$ appear more frequently, either explicitly or implicitly, in the domain-decontextualized annotations $\{e_{d}^{i}\}_{i=1}^{n_{d}}$. To ensure that explicitly mentioned features receive higher importance scores, we instruct $\mathcal{M}_{sel}$ to assign greater weight to features that are explicitly referenced in the annotations. Further details are in Appendix D.1.
Incremental Reinforcement Learning-based Feature Selection. When new data arrives, the offline LLM-based approach requires re-querying both old and new data, which becomes burdensome due to LLMs’ limited context window. As annotations increase, re-querying all data becomes impractical and costly, leading to higher resource consumption and reduced cost-effectiveness.
To address the limitations of the offline approach, we propose an Incremental Reinforcement Learning-based Feature Selection method that is more cost-effective for dynamic environments with evolving data. Specifically, we introduce a multi-agent reinforcement learning (MARL) framework to train two policy networks, $\mathcal{F}_{t}$ and $\mathcal{F}_{l}$, to select the top-$k$ most important time-series-wise and text-wise features, respectively. These policy networks store knowledge from existing annotations and are incrementally updated as new data arrives. This reduces the need to re-query the LLM with all the data, requiring only the new data during updates. As shown in Figure 1, each policy network is initialized with the first three layers of a small LLM, such as GPT-2 Radford et al. (2019), which remain frozen during training. A trainable multi-head attention layer and a language model (LM) head from GPT-2 follow these layers, using the smallest version of GPT-2 with 124M parameters.
@@
After training, the policy networks are incrementally updated with only new data, eliminating the need to re-query the LLM with both old and new data. This approach improves the scalability and efficiency of feature selection while reducing computational costs, effectively overcoming the offline approach’s limitations. By incrementally updating the policy networks, we ensure that feature selection remains scalable and cost-effective in dynamic environments with evolving data.
@@ 4.1 Experimental Setup
Dataset. To evaluate the effectiveness of TESSA, five real-world datasets from distinct domains are considered: Stock, Health, Energy, Environment, and Social Good. Specifically, the stock dataset includes 1,935 US stocks with the recent 6-year data, collected by ourselves. The other four datasets come from the public benchmark Time-MMD Liu et al. (2024a). In this paper, the Stock and Health datasets serve as the source domains, while the Energy, Environment, and Social Good datasets are treated as the target domains. Additionally, we generate a synthetic dataset containing both time series and ground-truth annotations to directly assess the quality of general annotations. More details on these datasets can be found in Appendix E.1. The rest four datasets are from a public benchmark Time-MMD. Additionally, we create a synthetic dataset with both time series and ground-truth annotations to directly evaluate the quality of general annotations. More details of these datasets are in Appendix E.1.
@@ 4.2 Evaluating General Annotations in Downstream Tasks
In this subsection, to evaluate the quality of the general annotations, we apply the generated annotations to the multi-modal downstream tasks (i.e., time series forecasting and imputation) by following the experimental setup in Time-MMD Liu et al. (2024a). The implementation details are provided in Appendix  F.1.
Baselines. To the best of our knowledge, TESSA is the first work to study cross-domain multi-modal time series annotation. To demonstrate its effectiveness, we compare it with several representative methods, including No-Text, Time-MMD Liu et al. (2024a), and DirectLLM as baselines. More details on these methods are in Appendix E.2.
Evaluation Metrics. For the time series forecasting task, we use MSE (Mean Squared Error) and MAE (Mean Absolute Error) as evaluation metrics, where lower values for both MSE and MAE mean better annotations.
Experimental Results. Table 1 presents the comparison results for the time series forecasting task, where Informer Zhou et al. (2021) is the forecasting model and GPT-4o Achiam et al. (2023) serves as the LLM backbone. Additional forecasting results using different LLM backbones are available in Appendix F.2. The following observations can be made: (1) No-Text shows the worst performance across all datasets, validating the need for annotations to improve performance in downstream tasks. This suggests that better downstream task performance indicates higher-quality annotations. (2) TESSA achieves the best performance among all compared methods, demonstrating its effectiveness in generating high-quality general annotations. Additional results of time series imputation tasks, can be found in Appendix F.3.

### v2

@@ Abstract
Time series data is ubiquitous across various domains, including manufacturing, finance, and healthcare. High-quality annotations are essential for effectively understanding time series and facilitating downstream tasks. However, obtaining such annotations is challenging, particularly in mission-critical domains. In this paper, we propose TESSA, a multi-agent system designed to automatically generate both general and domain-specific annotations for time series data. TESSA introduces two agents: a general annotation agent and a domain-specific annotation agent. The general agent captures common patterns and knowledge across multiple source domains, leveraging both time-series-wise and text-wise features to generate general annotations. Meanwhile, the domain-specific agent utilizes limited annotations from the target domain to learn domain-specific terminology and generate targeted annotations. Extensive experiments on multiple synthetic and real-world datasets demonstrate that TESSA effectively generates high-quality annotations, outperforming existing methods. Our code and data are available at https://anonymous.4open.science/r/TESSA-76C2.
@@ 1 Introduction
Time series data is prevalent in various fields such as manufacturing Hsu and Liu (2021), finance Lee et al. (2024), and healthcare Cascella et al. (2023). It captures critical temporal patterns essential for informed decision-making. However, general users frequently encounter difficulties in interpreting this data due to its inherent complexity, particularly in multivariate contexts where multiple variables interact over time. Furthermore, effective interpretation typically requires domain-specific knowledge to properly contextualize these patterns, thereby posing significant challenges for individuals without specialized expertise.
@@
High-quality annotations are crucial for addressing these interpretive challenges. Annotations provide meaningful context or insights into time series data, highlighting important patterns, events, or anomalies. They facilitate accurate analysis, forecasting, and decision-making, enhancing the performance of downstream tasks such as anomaly detection, trend prediction, and automated reporting. For instance, in predictive maintenance, understanding sensor data trends is vital for preventing equipment failure, while in finance, interpreting stock price movements is crucial for informed investment strategies. Despite their importance, high-quality annotations are often scarce in real-world applications. This scarcity stems primarily from the reliance on domain experts for manual annotation, which is resource-intensive, costly, and prone to inconsistencies. Moreover, the need for precise and domain-specific terminology further complicates the annotation process, as different fields require highly specialized knowledge for accurate and contextually relevant interpretation.
Figure 1: How to annotate time series across different domains?
@@
To alleviate the above issues, one straightforward approach is to leverage external resources to generate annotations Liu et al. (2024a). For example, Time-MMD Liu et al. (2024a) uses web searches to retrieve information as annotations, aiming to find similar patterns and descriptions from the internet. Others Jin et al. (2024); Liu et al. (2024b) directly adopt large language models (LLMs) for annotation, leveraging LLMs’ great language understanding capability. Prototype-based methods, such as prototype networks Ni et al. (2021), have also been employed to identify representative examples for annotation. However, these methods often fall short of producing high-quality annotations. Web search-based methods may retrieve irrelevant or inconsistent information. LLMs, while powerful, tend to generate generic annotations, capture only basic patterns, or even hallucinate, and fail to account for the complex nature of time series data. Prototype networks rely on large amounts of data to train the network and identify representative prototypes, but the scarcity of high-quality annotations limits the quality and representativeness of these prototypes, making it difficult to generalize effectively to new or unseen patterns.
@@
To address these limitations, we propose to extract knowledge from existing annotations across multiple source domains and transfer this knowledge to target domains with limited annotations. Specifically, as shown in Fig. 1, we aim to develop a system that automatically interprets time series data across various fields using common or domain-specific language. Formally, given abundant annotations from multiple source domains and limited annotations from a target domain, our goal is to leverage both time-series-wise and text-wise knowledge to generate accurate and contextually appropriate annotations for the target domain. There are two major technical challenges in developing such a system: (i) How to extract common knowledge from source domains? (ii) How to learn domain-specific jargon from limited target-domain annotations?
@@
To tackle these challenges and overcome the limitations of existing methods, we propose TESSA, a multi-agent system designed for both general and domain-specific TimE SerieS Annotation. As illustrated in Fig. 2, TESSA introduces two agents: a general annotation agent and a domain-specific annotation agent. The general annotation agent focuses on capturing common patterns and knowledge across various domains to generate annotations understandable by general users. To learn common knowledge from multiple domains, the general agent employs a time series-wise feature extractor and a text-wise feature extractor to extract both time-series-wise and text-wise features from time series data and domain-specific annotations from multiple source domains. To ensure important features are included in the general annotations, two feature selection methods—LLM-based and reinforcement learning-based selection—are introduced to effectively and efficiently select both the top-$k$ most important time-series-wise and text-wise features. The domain-specific agent leverages limited target-domain annotations to learn and generate annotations for specific domains using domain-specific terminologies (jargon). It incorporates a domain-specific term extractor to learn jargon from the limited target-domain annotations. Additionally, an annotation reviewer is proposed to maintain consistency between general annotations and domain-specific annotations.
@@
Our contributions are: (i) Problem. We explore a novel problem in cross-domain multi-modal time series annotation, bridging the gap between general understanding and domain-specific interpretation; (ii) Framework. We propose a novel multi-agent system, TESSA, designed for both general and domain-specific time series annotation by leveraging both time-series-wise and text-wise knowledge from multiple domains; (iii) Experiments. Extensive experiments on multiple synthetic and real-world datasets demonstrate the quality of the general and domain-specific annotations generated by TESSA.
@@ 2 Related Work
LLMs for Time Series Analysis. Recent advancements in LLMs have showcased their strong capabilities in sequential modeling and pattern recognition, opening up promising new directions for time series analysis. Several studies Xue and Salim (2023); Yu et al. (2023); Gruver et al. (2024); Jin et al. (2024); Li et al. (2024) have explored how LLMs can be effectively leveraged in this context. For instance, PromptCast Xue and Salim (2023) is a pioneering work that applies LLMs to general time series forecasting using a sentence-to-sentence approach. Yu et al. Yu et al. (2023) extend this by investigating the application of LLMs to domain-specific tasks, such as financial time series forecasting. LLMTime Gruver et al. (2024) demonstrates the efficacy of LLMs as time series learners by employing text-wise tokenization to represent time series data. Time-LLM Jin et al. (2024) reprograms time series data into textual prototypes for input into LLaMA-7B, enriched with natural language prompts that include domain expert knowledge and task-specific instructions. More details of related works are in Appendix A.
@@ 3.1 Multi-modal Feature Extraction
Time Series Feature Extraction. We extract features from time series data through a structured process $\mathcal{M}_{r}$. Formally, for each channel $c\in C$, the set of time-series features $\mathbf{F}_{t}$ is denoted as:
@@
where $\mathcal{M}_{r}$ denotes the feature extraction framework applied to $\mathbf{X}$, $f^{i}_{t}$ is the $i$-th extracted feature of $\mathbf{X}$, and $n_{t}$ is the number of extracted features. For multivariate time series data, inter-variable features (e.g., Pearson correlation) are also included. Details of $\mathcal{M}_{r}$ are provided in Appendix C.
@@
To address these challenges and facilitate knowledge transfer from source to target domains, we introduce a domain decontextualization LLM to convert domain-specific annotations into general annotations by removing domain-specific terminology. This makes it easier to extract common knowledge across domains. Specifically, given a domain-specific annotation $e_{s}^{i}$ in domain $d_{i}$, the decontextualized annotation $e_{d}^{i}$ is obtained as:
$e_{d}^{i}=\mathcal{M}_{d}(p_{de}(e_{s}^{i},d_{i})),$ (2)
@@
where $p_{l}$ is the prompt for text feature extraction to guide $\mathcal{M}_{l}$ to output the text-wise features explicitly or implicitly mentioned in the decontextualized annotations.
@@ 3.2 Adaptive Feature Selection
With a diverse set of features extracted from time series and text data, it becomes essential to focus on the most relevant ones to ensure the generated annotations remain concise and interpretable. Moreover, repeatedly querying LLMs with both the old and new data [^1] To avoid redundancy, unless specified otherwise, ‘data’ in this paper refers to time series and their corresponding textual annotations from various domains. each time wastes computational resources and incurs additional costs, especially when using non-open-source models.
@@
Offline LLM-based Feature Selection. Leveraging LLMs’ reasoning abilities, we introduce a feature selection method using LLM-generated feature importance scores to identify the top-$k$ most important time-series-wise and text-wise features. Features mentioned more frequently—either explicitly or implicitly—in annotations are assigned higher importance scores.
@@
Here, $p_{score}$ is the prompt used to score feature importance. Higher scores, $s_{j}$ and $s_{k}\in\mathbb{R}^{+}$, indicate that the features $f_{t}^{j}$ and $f_{l}^{k}$ appear more frequently, either explicitly or implicitly, in the domain-decontextualized annotations $\{e_{d}^{i}\}_{i=1}^{n_{d}}$. To ensure that explicitly mentioned features receive higher importance scores, we instruct $\mathcal{M}_{sel}$ to assign greater weight to features that are explicitly referenced in the annotations. More details are provided in Appendix D.1.
Incremental Reinforcement Learning-based Feature Selection. When new data11footnotemark: 1 arrives, the offline LLM-based approach requires re-querying both old and new data, which becomes burdensome due to LLMs’ limited context window. As annotations increase, re-querying all data becomes impractical and costly, leading to higher resource consumption and reduced cost-effectiveness.
To address the limitations of the offline approach, we propose an Incremental Reinforcement Learning-based Feature Selection method that is more cost-effective for dynamic environments with evolving data. Specifically, we introduce a multi-agent reinforcement learning (MARL) framework to train two policy networks, $\mathcal{F}_{t}$ and $\mathcal{F}_{l}$, to select the top-$k$ most important time-series-wise and text-wise features, respectively. These policy networks store knowledge from existing annotations and are incrementally updated as new data arrives. This reduces the need to re-query the LLM with all the data, requiring only the new data during updates. As shown in Fig. 2, each policy network is initialized with the first three layers of a small LLM, such as GPT-2 Radford et al. (2019), which remain frozen during training. A trainable multi-head attention layer and a language model (LM) head from GPT-2 follow these layers, using the smallest version of GPT-2 with 124M parameters.
@@
After training, the policy networks are incrementally updated with only new data, eliminating the need to re-query the LLM with both old and new data. This approach improves the scalability and efficiency of feature selection while reducing computational costs, effectively overcoming the offline approach’s limitations. By incrementally updating the policy networks, we ensure that feature selection remains scalable and cost-effective in dynamic environments with evolving data. More discussion of the necessity of the RL component is provided in Appendix D.2.
@@ 4.1 Experimental Setup
Dataset. To evaluate the effectiveness of TESSA, five real-world datasets from distinct domains are considered: Stock, Health, Energy, Environment, Social Good, Climate and Economy. Specifically, the stock dataset includes 1,935 US stocks with the recent 6-year data, collected by ourselves. The other four datasets come from the public benchmark Time-MMD Liu et al. (2024a). In this paper, the Stock and Health datasets serve as the source domains, while the rest five datasets are treated as the target domains. Additionally, we generate a synthetic dataset containing both time series and ground-truth annotations to directly assess the quality of general annotations. More details on these datasets can be found in Appendix F.1.
@@ 4.2 Evaluating General Annotations in Downstream Tasks
To evaluate the quality of the general annotations, we apply the generated annotations to the multi-modal downstream tasks (i.e., time series forecasting and imputation) by following the experimental setup in Time-MMD Liu et al. (2024a). As in Fig. 4, time series data and textual annotations are processed independently by unimodel TSF models and LLMs with projection layers. The model-specific outputs are then fused through a linear weighting mechanism to generate final predictions. Incremental RL-based selection in Section 3.2 is used in TESSA to select the top-$k$ most important features for generating annotations. The implementation details are provided in Appendix F.3 and G.1.
Baselines. TESSA is, to the best of our knowledge, the first work on cross-domain multi-modal time series annotation. We compare it with several representative single-domain methods: No-Text, Time-MMD Liu et al. (2024a), and DirectLLM (which directly uses LLM-generated annotations). Details of these methods are provided in Appendix F.2.
Evaluation Metrics. For time series forecasting task, we use MSE (Mean Squared Error) and MAE (Mean Absolute Error) as evaluation metrics, where lower values for both MSE and MAE mean better annotations.
Experimental Results. Table 1 presents the comparison results for the time series forecasting task, where Informer Zhou et al. (2021) is the forecasting model and GPT-4o Achiam et al. (2023) serves as the LLM backbone. Additional forecasting results using different LLM backbones are available in Appendix G.2. The following observations can be made: (1) No-Text shows the worst performance across all datasets, validating the need for annotations to improve performance in downstream tasks. This suggests that better downstream task performance indicates higher-quality annotations. (2) TESSA achieves the best performance among all compared methods, demonstrating its effectiveness in generating high-quality general annotations. Additional results of time series imputation tasks, can be found in Appendix G.3.